# The Dynamic, the Static, and the Weak factor models and the analysis of high-dimensional time series

### v1

@@ Abstract
Several fundamental and closely interconnected issues related to factor models are reviewed and discussed: dynamic versus static loadings, rate-strong versus rate-weak factors, the concept of weakly common component recently introduced by Gersing et al. (2023), the irrelevance of cross-sectional ordering and the assumption of cross-sectional exchangeability, and the problem of undetected strong factors.
@@ 1 Introduction: Spearman and the origins of factor models
Factor models and factor model methods are rooted in early-twentieth-century psychometrics. It is usually admitted that the concept of factor first appears more than a century ago in Spearman (1904) where a factor model is proposed in order to account for the dependencies between several variables related with cognitive abilities measured on given individuals. The result was a two-factor theory in which cognitive performance was explained by two unobservable “factors.” The objective of these early factor models, thus, is to account for cross-sectional dependencies—more specifically, cross-sectional covariances or correlations—in an i.i.d. context.
@@
    $\bm{\epsilon}_{t}=(\epsilon_{1t},\ldots,\epsilon_{nt})^{\prime}$, $t=1,\ldots,T$ is an unobserved sample from an i.i.d.  process of $n$-dimensional variables with ${\rm E}[{\bm{\epsilon}}_{t}]={\bm{0}}$ and ${\rm E}[{\bm{\epsilon}}_{t}{\bm{\epsilon}}_{t}^{\prime}]$ diagonal; call ${\bm{\xi}}_{t}=(\xi_{1t},\ldots,\xi_{nt})^{\prime}\coloneqq\bm{\epsilon}_{t}$ the idiosyncratic component of ${\bf X}_{t}$;
@@
Equation (1.1) thus decomposes the component $X_{it}$ of the observation ${\bf X}_{t}$ into two unobservable and mutually orthogonal components: the common component ${\chi}_{it}$ and the idiosyncratic component ${\xi}_{it}$. All subsequent factor models are based on such a decomposition, with various restrictions on $\chi_{it}$ and ${\xi}_{it}$: call it the factor model decomposition. Because the idiosyncratic $\bm{\xi}_{t}$ in (1.1) is a vector of $n$ mutually orthogonal white noises, this traditional model is called an exact factor model. The nature of (1.1), along with assumptions (a)–(e), is that of a semiparametric statistical model with parameters the loadings $\bf B$ and nuisances the unspecified distributions of ${\bf f}_{t}$ and $\bm{\epsilon}_{t}$.
@@
Estimation of this exact static factor model is typically performed via Gaussian MLE, which gives consistent estimates only for the loadings as $T\to\infty$, while the factors (which can be treated as incidental parameters) cannot be consistently estimated since $n$ is fixed (Anderson and Rubin, 1956; Lawley and Maxwell, 1971; Amemiya et al., 1987). The MLE, moreover, also requires finite 4th order moments of the observables.
@@
The observations in (1.1) being i.i.d., this exact model is of limited interest in the analysis of econometric data, where serial dependence is ubiquitous. The development of factor models in time-series and econometrics only began in the late seventies with the dynamic exact factor models of Geweke (1977) and Sargent and Sims (1977), who allow for factor-induced serial dependence and, a few years later, with the static approximate factor model of Chamberlain (1983) and Chamberlain and Rothschild (1983), who for the first time consider high-dimensional asymptotics (with both $n$ and $T$ tending to infinity) in the context.
@@
The Chamberlain and Rothschild approach has been popularized at the turn of the century thanks to the contributions of Stock and Watson (2002a, b), Bai and Ng (2002), Bai (2003), and many others. Few years earlier Forni et al. (2000) and Forni and Lippi (2001) had introduced the Generalized or General Dynamic Factor Model (GDFM) which nests all other factor models. The static model of Chamberlain and Rothschild, however, by far remain the most popular form of factor models in econometrics.
@@
In Section 2 we review the static approximate factor model of Chamberlain (1983) and Chamberlain and Rothschild (1983). Section 3 deals with the GDFM and, in particular, revisits Hallin and Lippi (2013) by showing that the GDFM decomposition is actually a representation result rather than just a statistical model. In Section 4 we reconsider the static approach of Chamberlain and Rothschild and show that, although less general, it can also be derived as a representation result; the proofs are developed in the Appendix. In Section 5 we discuss different concepts of weak factors, both static and dynamic. In 6 we show, both empirically and theoretically, that weak factors, in the usual acception of rate-weak factors, are incompatible with the a assumption of cross-sectional exchangeability. Section 7 discusses the crucial role of idiosyncratic components in relation to undetected factors. In Section 8 we conclude.
@@ 2 The Static: Chamberlain and Rothschild
    the $n\times 1$ observations ${\bf X}_{1},\ldots,{\bf X}_{T}$ and the $r\times 1$ factors ${\bf f}_{1},\ldots,{\bf f}_{T}$ are no longer i.i.d.;
@@
    the idiosyncratic components ${\bm{\xi}}_{t}^{\scriptscriptstyle\text{\rm{stat}}}=({\xi}_{1t}^{\scriptscriptstyle\text{\rm{stat}}},\ldots,{\xi}_{nt}^{\scriptscriptstyle\text{\rm{stat}}})^{\prime}$, $t=1,\ldots,T$ are no longer i.i.d. white noise, and
@@
Equation (2.1), thus, allows for serial dependencies among the $X_{it}$’s, hence for the analysisof $n$-dimensional time series.
Since ${\bm{\xi}}_{t}^{\scriptscriptstyle\text{\rm{stat}}}$ is no longer white noise and can be (serially and cross-sectionally) correlated, the model is called an approximate factor model (as opposed to the exact ones). As a consequence, it is not identified for fixed $n$, even under the classical identification constraints. This motivates another crucial novelty: high-dimensional asymptotics are considered, with both $n$ and $T$ going to infinity, where $n\to\infty$ yields asymptotic identification and $T\to\infty$ allows for consistent estimation as in the classical case. The loadings, however, remain static as in (1.1).
@@
    letting $\chi_{it}^{\scriptscriptstyle\text{\rm{stat}}}\coloneqq{\bf B}_{i}{\bf f}_{t}$, $\bm{\chi}_{t}^{\scriptscriptstyle\text{\rm{stat}}}\coloneqq(\chi_{1t}^{\scriptscriptstyle\text{\rm{stat}}},\ldots,\chi_{nt}^{\scriptscriptstyle\text{\rm{stat}}})^{\prime}$ (the statically common component) is the value at time $t$ of the unobserved finite-$(n,T)$ realization of a latent second-order stationary process
@@
    denoting by ${\bm{\Gamma}}^{(n)}_{{\chi}}$ its $n\times n$ lag-zero covariance matrix (which has rank $r$), the $r$ nonzero eigenvalues of ${\bm{\Gamma}}^{(n)}_{{\chi}}$ tend to infinity as $n\to\infty$;
@@
    with mean zero and $n\times n$ lag-zero covariance matrix ${\bm{\Gamma}}^{(n)}_{{\xi}}$; as $n\to\infty$, the eigenvalues of ${\bm{\Gamma}}^{(n)}_{{\xi}}$ remain bounded;
@@
    the factors and the idiosyncratic component (hence also the common and the idiosyncratic components) are mutually orthogonal: ${\rm E}[{f}_{kt}\,{\xi}_{it}]=0$, $k=1,\ldots,r$, $i\in\mathbb{N}$, and $t\in\mathbb{Z}.$
This static approximate factor model, like the exact model (1.1), constitutes a semiparametric statistical model, with parameters the loadings ${\bf B}_{i}$, $i=1,\ldots,n$ and nuisance the unspecified distributions of the idiosyncratic process $\bm{\xi}$ and the factors. Most people interpret the decomposition (2.1) (including the particular form of the common) as describing a data-generating process, which is irrelevant from a statistical perspective. Some (e.g. Onatski, 2012, Bai and Li, 2016) even treat the values of the factors as parameters.
Chamberlain and Rothschild do not consider the estimation problem for their approximate static factor model. Nor do they impose any rate on the divergence, as $n\to\infty$, of the eigenvalues of the covariance matrix ${\bm{\Gamma}}^{(n)}_{{\chi}}$, which in principle can be sublinear, linear, or superlinear.
Estimation is considered in Stock and Watson (2002a), Bai (2003), and Forni et al. (2009), among many others, who provide a rigorous treatment of the asymptotic properties of PCA-based estimators for the loadings and the factors of the Chamberlain and Rothschild model and show that, as expected, as both $n$ and $T$ tend to infinity, consistency (up to orthogonal transformations, as usual) is achieved for the factors and the loadings. Finite fourth-order moments of the observables are also required. QMLE methods also have been considered under the same set of assumptions (Doz et al., 2012; Bai and Li, 2016; Barigozzi and Luciani, 2024).
Typically, once factors are extracted (or estimated, with a slight abuse of language when they are not parameters) from an $n$-dimensional ($n$ large) time series ${\bf X}_{t}$ observed over $t=1,\ldots,T$, they are used in a second step to predict a given set of target variables (Stock and Watson, 2002a, b; Bai and Ng, 2006). This approach, in general, offers sizeable improvements over univariate or small-$n$ forecasting models and, to this day, this static approximate factor model remains one of the most popular tools in the analysis of high-dimensional time series.
@@ 3.1 A forerunner: the dynamic exact factor model of Geweke, Sargent, and Sims
Variants of this dynamic exact approach have been proposed by several authors, among which Engle and Watson (1981), Shumway and Stoffer (1982), Watson and Engle (1983), and Quah and Sargent (1993), who develop a “state-space approach” where dynamics are introduced by adding a parametric dynamic equation (typically, a VAR specification) for the factors to the static factor model decomposition (2.1). A high-dimensional generalization of the latter approach was proposed by Doz et al. (2011, 2012), but we do not discuss it here.
@@ 3.2 The General Dynamic Factor Model
The General Dynamic Factor Model (henceforth GDFM) proposed by Forni et al. (2000) and Forni and Lippi (2001) is combining the dynamic loadings ideas of the exact factor model of Geweke (1977) and Sargent and Sims (1977) with the flexibility of the approximate factor model of Chamberlain (1983) and Chamberlain and Rothschild (1983). The following presentation of the GDFM is inspired from the time-domain exposition of Hallin and Lippi (2013) and slightly differs from the original ones in Forni et al. (2000) and Forni and Lippi (2001). By avoiding the spectral-domain approach, it also avoids the assumption that a spectrum exists and the possibility of spectral eigenvalues diverging only on a subset of the frequency domain.
Differently from the previous ones in Sections 2 and 3.1, the approach adopted here is purely nonparametric (as opposed to semiparametric). The only fundamental assumption we need for the existence of the GDFM factor model decomposition (3.3) is that the observation, an $n\times T$ panel, is the finite realization, for $1\leq i\leq n$ and $1\leq t\leq T$ ($n$ and $T$ “large”), of some double-indexed second-order time-stationary stochastic process
@@
with (for convenience) mean zero: ${\rm E}[X_{it}]=0$, $i\in\mathbb{N}$, $t\in\mathbb{Z}$.
Instead of being part of the model specification and imposing constraints on the data-generating process, the factor model decomposition (3.3) (see below) into common and idiosyncratic is canonical and “endogenous” one, in the sense that any second-order time-stationary process of the form (3.2) determines such a decomposition; and, when additional conditions are imposed, they only involve the distribution of the observations $X_{it}$ (typically, their spectral density eigenvalues). This is in sharp contrast with the previous approaches by Chamberlain and Rothschild and Geweke, Sargent, and Sims, where the factor model decompositions (2.1) and (3.1) are part of the assumptions, along with conditions on their elements (typically, on the eigenvalues of the common and idiosyncratic covariance matrices ${\bm{\Gamma}}^{(n)}_{\chi}$ and ${\bm{\Gamma}}^{(n)}_{\xi}$).
@@
Call (3.3) the General Dynamic Factor (GDFM) decomposition of $\bf X$. Contrary to (1.1) and (2.1), this decomposition is not imposed on the data; it is canonical, always exists, and (beyond second-order stationarity) does not impose any restriction on the data-generating process of ${\bf X}$. In that sense, it is not a statistical model involving parameters, and is based on a canonical representation result. Whether that representation result is describing a data-generating process or not is irrelevant from a statistical perspective.
@@
$\theta\mapsto{\bm{\Sigma}}^{(n)}_{\bf X}({\theta})\coloneqq\sum_{k=-\infty}^{\infty}{\rm E}[{\bf X}_{t}^{(n)}{\bf X}_{{t-k}}^{(n)\prime}]{\exp}(-{\iota}k\theta),\quad\theta\in(-\pi,\pi]$
@@
for some collection ${\bf B}_{i}(L)\coloneqq\left(B_{i1}(L),\ldots,B_{iq}(L)\right)$ of linear square-summable filters $B_{ij}(L)$, $j=1,\ldots,q$ for all $i\in\mathbb{N}$; ${\bf u}_{t}$ then is the standardized innovations of ${\bm{\chi}}^{\scriptscriptstyle\text{\rm dyn}}$. These filters (for $i=1,\ldots,n$), moreover, can be expressed in terms of the eigenvalues and eigenvectors of ${\bm{\Sigma}}^{(n)}_{\bf X}({\theta})$. As a consequence, the GDFM decomposition (3.5) is identified.
@@
Under assumption (3.4), the dynamically common space $\mathcal{H}_{\scriptscriptstyle\text{\rm dyn com}}^{\bf X}$ of $\bf X$, moreover, is spanned by the values (at any time $t_{0}\in{\mathbb{Z}}$) of the $q$ first dynamic principal components $\{\psi_{jt}|\,t\in{\mathbb{Z}}\}$, $j=1,\ldots,q$ of $\bf X$ (see Brillinger (1981) or Hallin et al. (2018) for definitions). These dynamic principal components and their empirical counterparts play, in the GDFM and its estimation method as proposed in Forni et al. (2000), a role similar to that of traditional principal components in the static factor model (2.1) of Chamberlain (1983) and Chamberlain and Rothschild (1983) and its estimation as proposed in Stock and Watson (2002a), Bai (2003), etc. Being based on spectral density matrices, this technique requires finite moments of order four, the existence of a spectrum, and the existence of an integer $0\leq q<\infty$ such that (3.4) holds; note, however, that these assumptions, contrary to the assumptions (d) and (e) underlying (2.1), only involve the observable process $\bf X$. That technique, unfortunately, involves two-sided filters, hence performs poorly at the ends of the observation period—making it unsuitable in a forecasting context. An alternative estimation method involving only one-sided filters is studied in Forni et al. (2015), Forni et al. (2017), and Barigozzi et al. (2024).

### v3

@@ Abstract
Several fundamental and closely interconnected issues related to factor models are reviewed and discussed: dynamic versus static loadings, rate-strong versus rate-weak factors, the concept of weakly common component recently introduced by Gersing (2023), the irrelevance of cross-sectional ordering and the assumption of cross-sectional exchangeability, the impact of undetected strong factors, and the problem of combining common and idiosyncratic forecasts. Conclusions all point to the advantages of the General Dynamic Factor Model approach of Forni et al. (2000) over the widely used Static Approximate Factor Model introduced by Chamberlain and Rothschild (1983).
@@ 1 Introduction: Spearman and the origins of factor models
Factor models and factor model methods are rooted in early-twentieth-century psychometrics. It is usually admitted that the concept of factor first appears, more than a century ago, in Spearman (1904) where a factor model is proposed in order to account for the dependencies between several variables related with cognitive abilities measured on given individuals. The result was a two-factor theory in which cognitive performance was explained by two unobservable “factors.” The objective of these early factor models, thus, is to account for cross-sectional dependencies—more specifically, cross-sectional covariances or correlations—in an i.i.d. context.
@@
    $\bm{\epsilon}_{t}=(\epsilon_{1t},\ldots,\epsilon_{nt})^{\prime}$, $t=1,\ldots,T$ is the unobserved realization of an i.i.d.  process of $n$-dimensional variables with ${\rm E}[{\bm{\epsilon}}_{t}]={\bm{0}}$ and ${\rm E}[{\bm{\epsilon}}_{t}{\bm{\epsilon}}_{t}^{\prime}]$ diagonal; call ${\bm{\xi}}_{t}=(\xi_{1t},\ldots,\xi_{nt})^{\prime}\coloneqq\bm{\epsilon}_{t}$ the idiosyncratic component of ${\bf X}_{t}$;
@@
Equation (1.1) thus decomposes the component $X_{it}$ of the observation ${\bf X}_{t}$ into two unobservable and mutually orthogonal components: the common component ${\chi}_{it}$ and the idiosyncratic component ${\xi}_{it}$. All subsequent factor models are based on such a decomposition, with various restrictions on $\chi_{it}$ and ${\xi}_{it}$: call it the factor model decomposition. Because the idiosyncratic $\bm{\xi}_{t}$ in (1.1) is a vector of $n$ mutually orthogonal white noises, this traditional model is called an exact factor model. The nature of (1.1), along with assumptions (a)–(e), is that of a semiparametric statistical model with parameters the loadings $\bf B$ and nuisances the unspecified joint distribution of ${\bf f}_{t}$ and $\bm{\epsilon}_{t}$. Note that, because of i.i.d-ness in (c) and (d), the orthogonality condition in (e) holds between ${f}_{kt}$ and ${\epsilon}_{it^{\prime}}$ for all $k=1,\ldots,r$, $i=1,\ldots,n$, and $t=1,\ldots,T$, but also for all $t^{\prime}=1,\ldots,T$. If $t$ is interpreted as time, this means that (e) implies orthogonality between $\bm{\chi}_{t}$ and $\bm{\xi}_{t}$ at all leads and lags; the same implication no longer holds in subsequent sections, where $\bm{\chi}_{t}$ and $\bm{\xi}_{t}$ are time series exhibiting serial dependence.
@@
Estimation of this exact static factor model is typically performed via Gaussian MLE, which gives consistent (as $T\to\infty$) estimates only for the loadings, while the factors (which can be treated as incidental parameters) cannot be consistently estimated since $n$ is fixed (Anderson and Rubin, 1956; Lawley and Maxwell, 1971; Amemiya et al., 1987). The MLE, moreover, also requires finite fourth-order moments of the observables.
@@
The observations in (1.1) being i.i.d., this exact model is of limited interest in the analysis of econometric data, where serial dependence is ubiquitous. The development of factor models in time-series and econometrics only started in the late seventies with the dynamic exact factor models of Geweke (1977) and Sargent and Sims (1977), who allow for factor-induced serial dependence and, a few years later, with the static approximate factor model of Chamberlain (1983) and Chamberlain and Rothschild (1983), who for the first time consider high-dimensional asymptotics (with the dimension $n$ tending to infinity) in the context.
@@
The Chamberlain and Rothschild approach has been popularized at the turn of the century thanks to the contributions of Stock and Watson (2002a, b), Bai and Ng (2002), and Bai (2003), followed by many others. A few years earlier, Forni et al. (2000) and Forni and Lippi (2001) had introduced the Generalized or General Dynamic Factor Model (GDFM) which nests all other factor models. The static approximate model of Chamberlain and Rothschild, however, by far remains (with both $n$ and $T$ tending to infinity) the most popular and widespread form of factor model in applied econometrics.
@@
In Section 2, we briefly review the static approximate factor model of Chamberlain (1983) and Chamberlain and Rothschild (1983). Section 3 deals with the GDFM and, in particular, revisits Hallin and Lippi (2013) by showing that the GDFM factor model decomposition is actually a representation result rather than just a statistical model. In Section 5 we reconsider the static approach of Chamberlain and Rothschild and show that, although less general than the GDFM, it can also be derived as a representation result; the proofs are developed in the Appendix. Section 6 discusses different concepts of weak factors, both static and dynamic. In Section 7, we show, both empirically and theoretically, that weak factors, in the usual acception of rate-weak factors, are incompatible with the a assumption of cross-sectional exchangeability. Section 9 examines the crucial role of idiosyncratic components in relation to undetected factors, and the importance, in forecasting problems, of orthogonality at all leads and lags between the common and idiosyncratic components. Section 10 concludes.
@@ 2 The Static: Chamberlain and Rothschild
    the $n\times 1$ observations ${\bf X}_{1},\ldots,{\bf X}_{T}$ and the $r\times 1$ factors ${\bf f}_{1},\ldots,{\bf f}_{T}$ no longer are assumed to be i.i.d.,
@@
    the idiosyncratic components ${\bm{\xi}}_{t}^{\scriptscriptstyle\text{\rm{stat}}}=({\xi}_{1t}^{\scriptscriptstyle\text{\rm{stat}}},\ldots,{\xi}_{nt}^{\scriptscriptstyle\text{\rm{stat}}})^{\prime}$, $t=1,\ldots,T$ no longer are assumed to be  i.i.d. white noise, and
@@
Equation (2.1), thus, allows for serial cross-dependencies among the $X_{it}$’s, hence for the analysisof $n$-dimensional time series.
Since ${\bm{\xi}}_{t}^{\scriptscriptstyle\text{\rm{stat}}}$ is no longer white noise and can be (serially and cross-sectionally) correlated, the model is called an approximate factor model (as opposed to the exact ones). As a consequence, it is not identified for fixed $n$, even under the classical identification constraints. This motivates another crucial novelty: high-dimensional asymptotics are considered, with both $n$ and $T$ going to infinity, where $n\to\infty$ yields asymptotic identification while $T\to\infty$ allows for consistent estimation as in the classical case. The loadings, however, remain static as in (1.1)—whence the terminology static approximate factor model.
@@
    (d1) letting $\chi_{it}^{\scriptscriptstyle\text{\rm{stat}}}\coloneqq{\bf B}_{i}{\bf f}_{t}$, the $n$-dimensional vector $\bm{\chi}_{t}^{\scriptscriptstyle\text{\rm{stat}}}\coloneqq(\chi_{1t}^{\scriptscriptstyle\text{\rm{stat}}},\ldots,\chi_{nt}^{\scriptscriptstyle\text{\rm{stat}}})^{\prime}$ (the statically common component) is the value at time $t$ of the unobserved finite-$(n,T)$ realization of a latent second-order stationary process
@@
    (d2) denoting by ${\bm{\Gamma}}^{(n)}_{{\chi}}$ its $n\times n$ lag-zero covariance matrix (which has rank $r$), the $r$ nonzero eigenvalues of ${\bm{\Gamma}}^{(n)}_{{\chi}}$ tend to infinity as $n\to\infty$;
@@
    with mean zero and $n\times n$ lag-zero covariance matrix ${\bm{\Gamma}}^{(n)}_{{\xi}}$; as $n\to\infty$, all eigenvalues of ${\bm{\Gamma}}^{(n)}_{{\xi}}$ remain bounded;
@@
    the factors ${\bf f}_{t}$ and the idiosyncratic components $\xi_{it}$ (hence also the common and the idiosyncratic components) are mutually orthogonal at time $t$: ${\rm E}[{f}_{kt}\,{\xi}_{it}]=0$, $k=1,\ldots,r$, $i\in\mathbb{N}$, and $t\in\mathbb{Z}.$
This static approximate factor model, like the exact model (1.1), constitutes a semiparametric statistical model, with parameters the loadings ${\bf B}_{i}$, $i=1,\ldots,n$ and nuisance the unspecified distributions of the idiosyncratic process $\bm{\xi}$ and the factors. Most people interpret the decomposition (2.1) (including the particular form of the common) as describing a data-generating process, which is irrelevant from a statistical perspective. Some (e.g. Onatski, 2012, Bai and Li, 2016) even treat the values of the factors as parameters—i.e., condition their analysis on the unobserved latent factor process.
Chamberlain and Rothschild, however, do not include (d2) in their assumptions (see their Definition 2); as a result, the number of factors, in their version of (2.1), is not well identified (that number could be any $K\geq r$) [^1] We thank Marco Lippi for attracting our attention to this fact. and their factors need not be “pervasive” in the sense of inducing an unbounded (as $n\to\infty$) ${\bm{\Gamma}}^{(n)}_{{\chi}}$ eigenvalue. Nor do they impose any rate on the divergence, as $n\to\infty$, of the diverging eigenvalues of the covariance matrix ${\bm{\Gamma}}^{(n)}_{{\chi}}$, which in principle can be sublinear, linear, or superlinear. And they do not consider the estimation problem for their static approximate factor model.
Estimation is considered in Stock and Watson (2002a), Bai (2003), and Forni et al. (2009), among many others, who provide a rigorous treatment of the asymptotic properties of PCA-based estimators for the loadings and the factors of the Chamberlain and Rothschild model and show that, as expected, consistency (up to orthogonal transformations, as usual) is achieved, as both $n$ and $T$ tend to infinity, for the factors and the loadings. Finite fourth-order moments of the observables are also required. QMLE methods also have been considered under the same set of assumptions (Doz et al., 2012; Bai and Li, 2016; Barigozzi and Luciani, 2024).
Typically, once factors are extracted (or estimated, with a slight abuse of language when they are not considered as parameters) from an $n$-dimensional ($n$ large) time series ${\bf X}_{t}^{(n)}$ observed over $t=1,\ldots,T$, they are used in a second step to predict a given set of target variables (Stock and Watson, 2002a, b; Bai and Ng, 2006). This approach, in general, offers sizeable improvements over univariate or low-dimensional (small-$n$) forecasting methods and, to this day, this static approximate factor model –with (see Section 4) some variations on the orthogonality between common and idiosyncratic components—remains the most popular tool in the analysis of high-dimensional time series.
@@ 3.1 A forerunner: the dynamic exact factor model of Geweke, Sargent, and Sims
Variants of this dynamic exact approach have been proposed by several authors, among which Engle and Watson (1981), Shumway and Stoffer (1982), Watson and Engle (1983), and Quah and Sargent (1993), who develop a “state-space approach” where dynamics are introduced by adding to the static factor model decomposition (2.1) a parametric dynamic equation (typically, a VAR specification) for the factors. A high-dimensional generalization of the latter approach was proposed by Doz et al. (2011, 2012), which we do not discuss here.
@@ 3.2 The General Dynamic Factor Model
The General Dynamic Factor Model (henceforth GDFM) proposed by Forni et al. (2000) and Forni and Lippi (2001) is combining the dynamic loading features of the exact factor model of Geweke (1977) and Sargent and Sims (1977) with the flexibility of the approximate factor model of Chamberlain (1983) and Chamberlain and Rothschild (1983). The following presentation of the GDFM is inspired from the time-domain exposition of Hallin and Lippi (2013) and slightly differs from the original ones in Forni et al. (2000) and Forni and Lippi (2001). By avoiding the spectral-domain approach of Forni et al. (2000) and Forni and Lippi (2001), it also avoids the assumption that a spectrum exists and the possibility of spectral eigenvalues diverging only on a subset of the frequency domain.
Differently from the previous ones in Sections 2 and 3.1, the approach adopted here is purely nonparametric (as opposed to semiparametric). The only fundamental assumption required for the existence of the GDFM factor model decomposition (3.3) is that the observation, an $n\times T$ panel, is the finite realization, for $1\leq i\leq n$ and $1\leq t\leq T$ ($n$ and $T$ “large”), of some double-indexed second-order time-stationary stochastic process
@@
with (for convenience and without loss of generality) mean zero: ${\rm E}[X_{it}]=0$, $i\in\mathbb{N}$, $t\in\mathbb{Z}$.
Instead of being part of the model specification and imposing constraints on the data-generating process, the factor model decomposition (3.3) (see below) into common and idiosyncratic is canonical and “endogenous,” in the sense that any second-order time-stationary process of the form (3.2) admits such a decomposition; and, when additional conditions are imposed, they only involve the distribution of the observations $X_{it}$ (typically, their spectral density eigenvalues). This is in sharp contrast with the previous approaches by Chamberlain and Rothschild and Geweke, Sargent, and Sims, where the factor model decompositions (2.1) and (3.1) are part of the assumptions, along with conditions on their elements (typically, on the eigenvalues of the common and idiosyncratic covariance matrices ${\bm{\Gamma}}^{(n)}_{\chi}$ and ${\bm{\Gamma}}^{(n)}_{\xi}$).
@@
Call (3.3) the General Dynamic Factor Model (GDFM) decomposition of $\bf X$. Contrary to (1.1) and (2.1), this decomposition is not imposed on the data; it is canonical, always exists, and (beyond second-order stationarity) does not put any restriction on the data-generating process of ${\bf X}$. In that sense, it is not a statistical model involving parameters, and constitutes a canonical representation result. Whether that representation result is describing a data-generating process or not is irrelevant from a statistical perspective.
@@
$\theta\mapsto{\bm{\Sigma}}^{(n)}_{X}({\theta})\coloneqq\sum_{k=-\infty}^{\infty}{\rm E}[{\bf X}_{t}^{(n)}{\bf X}_{{t-k}}^{(n)\prime}]{\exp}(-{\iota}k\theta),\quad\theta\in(-\pi,\pi]$
@@
for some collection ${\bf B}_{i}(L)\coloneqq\left(B_{i1}(L),\ldots,B_{iq}(L)\right)$ of linear square-summable filters $B_{ij}(L)$, $j=1,\ldots,q$ for all $i\in\mathbb{N}$, with $\{{\bf u}_{t}|\,t\in{\mathbb{Z}}\}$ the standardized innovations of ${\bm{\chi}}^{\scriptscriptstyle\text{\rm dyn}}$. These filters (for $i=1,\ldots,n$) can be expressed in terms of the eigenvalues and eigenvectors of ${\bm{\Sigma}}^{(n)}_{X}({\theta})$. As a consequence, the GDFM decomposition (3.5) is identified.
@@
Under assumption (3.4), the dynamically common space $\mathcal{H}_{\scriptscriptstyle\text{\rm dyn com}}^{\bf X}$ of $\bf X$, moreover, is spanned by the values (at any time $t_{0}\in{\mathbb{Z}}$) of the $q$ first dynamic principal components $\{\psi_{jt}|\,t\in{\mathbb{Z}}\}$, $j=1,\ldots,q$ of $\bf X$ (see Brillinger (1981) or Hallin et al. (2018) for definitions). These dynamic principal components and their empirical counterparts play, in the GDFM and its estimation method as proposed in Forni et al. (2000), a role similar to that of traditional principal components in the static factor model (2.1) of Chamberlain (1983) and Chamberlain and Rothschild (1983) and its estimation as proposed in Stock and Watson (2002a), Bai (2003), etc. Being based on spectral density matrices, this technique requires finite moments of order four, the existence of a spectrum, and the existence of an integer $0\leq q<\infty$ such that (3.4) holds; note, however, that these assumptions, contrary to the assumptions (d) and (e) underlying (2.1), only involve the observable process $\bf X$. That technique, unfortunately, involves two-sided filters, hence performs poorly at the ends of the observation period—making it unsuitable in a forecasting context. An alternative estimation method involving only one-sided filters is proposed in Forni et al. (2015) and studied in Forni et al. (2017) and Barigozzi et al. (2024).